CH/Meet/93/01

Minutes

Charissa Group Meeting

Meeting held on: 10 February 1993

Meeting held at: Birmingham

Present:

BRF(Chair),WNC,NMC,GJG,SJAH,CDJ,ACM, VP,WDMR,GT,RPW,DLW,JZ

Apologies:  RAC,RAH,JSL
Distribution: Those present, plus RAC,RAH,JSL

Minutes

1. Minutes of Meeting of 25 November 1992: accepted

2. Matters Arising: actions were reviewed and found to be adequately covered. Item 8 — Charissa labels for electronics still to be pursued. DLW Item 9 — Noise in hybrid detector 3 may be in the distribution box (now in York). DLW suggested further tests. DLW Item 9 — Window thicknesses for ANU should be advised to RAC.

Item 10 — ANU data acquisition backup to be pursued.  WDMR,RAH
Item 14 — Student Training: DLW to prepare document.  DLW

3. Update on NSC/SERC matters: The meeting heard that the Nuclear Physics Board had forwarded advice to SERC regarding the implementation of a 2/4/6% progressive cut in order to create a flexibility fund for ABRC. It was understood that Nuclear Structure funding was entirely within the SERC bid for flexibility funds.

4. Progress on Detector Tests:

DLW distributed a document produced by S.P. Chappell (Ch/Tech/93/03) on Hamamatsu data taken at the end of the last run. Preliminary figures indicate excellent position resolution with minimal distortion. A computer version of

the document should be forwarded to WNC.  DLW,SPC

The sample detector from Intertechnique has still not arrived, but BRF had been advised that it was now in the UK.

The detectors from Intertechnique, Micron (NASA style detector with uncertain connections) and Hamamatsu (recently arrived strip detector similar to Micron 16–strip design) each require beam tests, which should be parasited

onto the next run.  DLW

NMC reported that John Caunt Scientific distribute position sensitive detectors made by Quadrant and have offered samples. It was agreed that these

should be investigated.  NMC

NMC showed a new detector testing box that allowed a simple test of continuity for strip connections in the 16–strip detectors.

5. Progress on Scintillators:

NMC reported that all 50 scintillators had been delivered on time to Birmingham, and had been tested with alpha-particles. He showed data comparing their α–particle response to the manufacturers’ specification for γ–ray response. All detectors worked and 80% exceed the prototype’s performance.

All detectors easily satisfy the size specification. NMC reported that the metal support clamps in the new modules would be replaced by teflon.

6. Progress on Chamber and Detector Assemblies: Regarding the chamber, DLW reported that the target changing assembly work could commence. It was, however, believed to be more appropriate to direct effort to detectors at this stage, until the chamber design has advanced further.

Regarding detectors, DLW reported that a prototype module had been produced at York.

Two backplates for the new modules will be built in York: one in aluminium, for Oxford to investigate preamp mounting, one in stainless, for full assembly of a module at York. The latest design features a new layout for the feedthroughs, to allow a tube to be brought out through the central position in each module.

Each module will include 10 ribbon cable vacuum feedthroughs. Enough for 6 modules have been ordered through DL; the plan is to make 5 modules initially, and to hold spares.

Completion of a prototype module is expected by the end of March.

7. Progress on Gas Supply:

DLW reported in place of RAC. A prototype will be mechanically complete by mid–March and hopefully tested by the end of March.

The control of the needle valve position had been changed back to conventional manual control, making a significant saving.

8. Progress on Electronics and Data Acquisition: WDMR reported for RAH, due to a mix-up in transport arrangements.

Preamplifiers: design essentially complete; awaiting final design for backplate.

NMC reported that the preamplifiers modified to high gain for gas detectors operated satisfactorily in the recent run, apart from some apparent reliability problems. One unit in each dual module had failed immediately. WDMR suggested that the actual modifications, rather than the design, was probably at fault. The failed boards should be removed and sent to Oxford for inspection.

NMC

Bias Supplies: prototypes are built, and P. Simmons is working on the microprocessor control program for compensation. NMC queried whether bias supplies for gas detectors were included in the plans. WDMR replied that tentatively there were 120 supplies planned, i.e. enough for gas biases also, but that the final number could not be confirmed until a full costing was completed.

Amplifiers/Gated Integrators: design is progressing, but currently not as top priority. Prototype expected by Easter, to allow tests with Ge detectors.

Manufacture: it is expected that much of this will need to be contracted out, especially the preamplifiers. Oxford and York to liaise on electronics effort.

WDMR,DLW

Non-Linearities: initial studies of a single Silena, by P. Simmons at Oxford, suggest the amplifier or preamp is to blame. This must be clarified before preamp production begins. WNC circulated results from a preliminary analysis by S.M. Singer of matchstick data taken by NMC on the recent run. Nonlinearities were evident.

WDMR described linearity tests required at the time of the next run. Tests without preamps, and with non-Charissa amplifiers should be included, and

the results sent to Oxford.  RPW

New Data Acquisition System: the 32 channel system along the lines of Phase 2 (cf. Ch/Tech/92/06) should be ready by the end of Summer 1993. Hence,

prototype testing in Canberra would be possible in March 1994.  RAH

9. Developments on Hybrid Counters: The performance in experiments was discussed with reference to a document written by A.St.J. Murphy (Ch/Tech/93/01). This showed that the position resolution was excellent (individual 0.1mm wires clearly seen) as was the energy resolution for a single strip (250 keV).

BRF pointed out that the optical transparency of the grids (90% per grid, 80% for the two grids) implies a significant loss in coincidence efficiency. The possibilities for modifying or removing grids should therefore be investigated.

RAC

WDMR suggested that this documented performance be used as a standard to judge the alternatives being assessed for the new modules. This was agreed.

Radiation-damaged strip detectors: the leakage for the central strips is increased relative to edge strips, but there is no explanation of why this should be.

10. Developments on Event Manager and GEC Emulator: Completion of Phase 0: VP reported that the emulator had been used for several hours during the run in January, and data had been written to exabyte for a two hour test period. Thus, Phase 0 is now working, and will be used for the major part of the data taking in the February experiment at the NSF. (For a discussion of the phases identified earlier, see Ch/Tech/92/06.)

The data transfer from CAMAC to VME is limited by the CAMAC cycle time in Phase 0. The VME processor is tied up dealing with this, which will result in a low sort fraction on-line. A system had been transported to Birmingham for tests.

Phase 1 Progress: In this Phase, a Starburst in CAMAC puts the event controller data into the data stack, reads the TDC’s and loads TDC data into the data stack. It then constructs the header and hit words and loads these into the data stack. Finally, the readout of the data stack into VME memory over the FERA/VME link is initiated. This will greatly speed the readout time and solve the problems causing a low sort fraction in Phase 0. Checks of this readout process are in preparation at Oxford. Most of the remaining work is in software, and P. Simmons is working on this. It is planned to use a PC to interogate the CAMAC and VME crates.

It was agreed that the higher sort fraction made it desirable to bring Phase 1 forward for use at the ANU in September if possible. Bench tests should be complete by the end of Summer. WDMR,RAH A serial line to Starburst and a cross assembler should be supplied to Oxford

to enable progress in Phase 1.  VP

11. Documentation of Experimental Setup:

BRF reviewed the agreed areas of responsibility, viz. Chamber and Gas (York), Electronics (Oxford), Data Acquisition (Birmingham), Running check list (Surrey), GEC guide (Birmingham). Progress on this had been understandably slow, but should pick up.

GJG circulated a draft checklist for run-time jobs, for comment.

WDMR reminded the meeting of an earlier document that he had produced, which included relevant details. BRF agreed to send a copy to GJG. BRF

12. Completion of Programme at NSF:

The next run was discussed (Proposal 900, due to start 19 Feb). It was possible that Proposal 904 (angular distribution) may also get time in the final schedule.

Orsay run: DLW reported that this experiment will not be able to go ahead in April/May due to the schedule for reconfiguring the chateau de cristal. An October/November timeslot is planned for the 21 shifts allocated.

13. Planning for ANU Experiment: The precise date of the run planned for the Southern Spring was re-opened for discussion, in view of the wish to avoid a clash with the Postgraduate Summer School at St. Andrews. Running at the end of September and early October received most support. It was thought that some students (not directly using the data from this run) could be involved in setting up, prior to the Summer School, but might miss the actual data collection in order to return for the Summer School.

Collaboration members should advise BRF of their availability and Term

times, and he will collate.  ALL

Shipping of equipment: DLW reported that an estimate from the Linac shippers had been obtained for 6 racks of electronics, one gas handling rig, large boxes of detectors etc., and two Sun workstations. Including documentation, customs and delivery to the ANU, this came to £4.5K with 10 weeks delivery. Insurance is extra.

DLW reported news from Dr D.C. Weisser (ANU), via Paddy Regan, that the ANU would be willing to cooperate in sharing space in a container. They would benefit from shipping some items in April rather than at the end of June when the Linac container is going.

Cable runs: GJG and WNC presented a document detailing the options for providing cable runs, and including specifications for various cables considered (Ch/Tech/93/02). It was agreed that the preferred option was to obtain superscreen cables from DL, that could be removed and re-used un-cut. These had been bid for as “minor equipment”. If this was unsuccessful, then ready-made cable should be purchased in Australia. The specification for superscreen cable should be faxed to an Australian

supplier to see whether an equivalent is available.  GJG

It was understood that installation was going to be time consuming and would require some coordination of effort.

Cables from feedthroughs in the vacuum chamber would be brought out to a patch panel located as near as possible to the chamber. The cables should be arranged in groups, brought to pads with an array of slide-on Sealectro female connectors, for connection to the vacuum flanges.

Vacuum feedthroughs: it was agreed that two flanges on the base of the 2metre chamber, presently used for BNC feedthroughs, should be replaced with flanges holding 60 Sealectro (conhex) feedthroughs each (total of 120). The hermetic feedthroughs should be snap-on to snap-on male fittings, arranged in arrays to mate with the pads of the cable groups.

Inside of chamber: the carriages used on the movable arms are obsolete. It was agreed to purchase the remaining stock of the UK supplier, viz. 3 rails

and 6 carriages, for £1K approximately.  DLW

Cables inside the chamber will be provided by York. These should mate to

the flanges using pads as described above.  DLW

Spacers to bring detectors to beam height are anticipated to be built in time

for shipping in April.  DLW

Computing: VP reported that three internet addresses were required. BRF VP advised that a hub unit to join our twisted pair connections to the ANU ethernet will need to be purchased. He also advised that the ANU might reasonably be expected to install the actual ethernet connection and necesssary transceiver.

First setting up: shipping times imply that equipment will start arriving at the end of June. BRF and SJAH were available for unpacking the first equipment and initial setting up at the beginning of July. Extra coverage would be needed in July and August.

Subsistence costs: the most satisfactory solution was thought to consist of an agreed per diem for subsistence being paid to each person, in addition to accommodation costs as incurred. Accommodation would be coordinated and was anticipated to involve a share house, with hostel and/or motel rooms to supplement at peak times.

BRF agreed to investigate bookings on the first visit.  BRF

Transport: some car hire was likely to be necessary. Car purchase was considered impractical. Members were invited to submit bicycles for shipping to Canberra, if there was space in the container.

SERC students: BRF reminded the meeting that he had written to the NSC secretary in November 1992, advising of the probable requirements for students visiting Australia in order to participate in Charissa experiments. He received an assurance that everything was in hand for providing suitable support when it was requested. It was decided that such requests should use a standard form of words agreed by the collaboration. Meanwhile, DLW pro-

posed that he would send in a trial form and this was agreed.  DLW

14. Plans for Strasbourg: BRF reported that JSL had discussed charged particle collaborations with Francis Beck during the latter’s visit to DL in January.

BRF produced a message from Dick Freeman (Strasbourg) regarding a meeting proposed for May 1993 in Strasbourg, to discuss possible future collaborative research with the Vivitron facility. It was agreed that a strong presence from the Charissa project would be desirable.

It was confirmed that the Charissa chamber and beamline from DL should be offered for siting at the Strasbourg laboratory, as a contribution to future collaborative efforts.

15. Legnaro Meeting:

DLW reported on the December 1992 meeting at Legnaro, Italy, to discuss European collaboration in nuclear reaction mechanism studies. From Charissa, DLW and JSL had attended. DLW circulated a summary of the meeting produced by Eric Plagnol (GANIL).

The research interests represented at the meeting ranged from sub-Coulomb barrier to GeV/A collisions. The most fruitful discussion centred on shared interests in instrumentation.

A useful contact was established with representatives of the MULTICS collaboration. Their equipment was being returned from a run at Michigan State for setting up at Catania, to await beam there.

DLW had agreed to organise, with A. Moroni (Milan), a working group meet-

ing on Silicon–CsI Telescopes during 1993.  DLW

16. Minor Equipment Bids: A bid for equipment needed for Charissa experiments abroad had been submitted from Birmingham, on behalf of the collaboration. Individual groups had bid for equipment to be located at their institutions, for use on Charissa developments, as a part of their own bids.

A copy of the bid should be circulated to each group.  BRF

Any ommissions should be notified to BRF as soon as possible.

WNC,NMC,WDMR,GT,DLW

Daresbury should be notified of any Charissa modules presently included in

the Eurogam set-up.  BRF

17. Glasgow Contributions: Six abstracts were planned for submission covering Charissa experiments: J. Pople, A. Murphy and M.Leddy (Birmingham), S. Fox (York), P. Simmons (Oxford) and S. Singer (Surrey).

18. Any Other Business: M. Freer: Martin Freer has succeeded in being awarded an SERC Advanced Fellowship for work with the Charissa group, based at Birmingham. He is expected to commence in Summer 1993.

Micron Strip Detectors: BRF reported that Micron had agreed to repair 10 damaged detectors for free. However, it was now clear that strip contacts were being damaged by our own handling. In future, Micron aim to use thicker PCBs. Meanwhile, increased care should be employed in handling, and all

group members should be made aware of this.  ALL

SERC Annual Report: this features a description of Charissa research (written by BRF).

Invited Talks: BRF would be talking at the Daresbury meeting on physics achievements at the NSF (March 1993).

BRF and ACM would talk at the International Conference on Atomic and Nuclear Clusters, Santorini (June 1993) [WDMR was also invited but cannot attend], WDMR will talk at the Glasgow IOP Conference (March 1993).

Unix Operating Systems: WNC requested that a clear strategy on system upgrades be developed, citing the Solaris 2.1 implementation as an example. (At Surrey, the Nuclear computers are integrated into the Departmental unix system.) It was agreed that DL would lead on implementing system upgrades, with Birmingham undertaking to upgrade collaboration software at an early

stage in each system change.  SJAH

Solaris 2.1 would not be implemented before the end of June 1993.

19. Date of Next Meeting: Maundy Thursday, 8 April 1993, Daresbury Laboratory, 11 a.m.

A room to be booked, and taxi arrangements clarified.  VP

W.N. Catford, Secretary. February 18, 1993 (released February 24, 1993)